# A Concept-Centric Approach to Multi-Modality Learning

### v1

@@ Appendix B Evaluation Datasets and Preprocessing
Appendix C Projection Models Details
C.1 Architecture
ViT-based vision-modality projection models use a vision transformer (ViT-Base) pretrained on ImageNet-21k Dosovitskiy et al. (2020) as the backbone. The baseline MLP model is comprised of three fully-connected layers used as ViT’s classification head, with each middle layer containing 128 neurons.
ResNet-based vision-modality projection models use a ResNet model (ResNet-50) pretrained on ImageNet-21k He et al. (2015) as the backbone. Because of ResNet’s large feature vectors, the linear layer used to project feature vectors onto the concept space is expanded to a three-layer MLP, featuring two intermediate layers comprising 512 and 256 neurons, respectively. The baseline MLP model is comprised of three fully-connected layers installed after ResNet’s layers, with each middle layer containing 128 neurons.
BERT-based nlp-modality projection models use a pretrained BERT encoder (BERT-base) Devlin et al. (2018) as the backbone.
C.2 Training Details
@@
Appendix D Image-Text Matching Experiment Details
D.1 Our Framework
We follow the cross-modality joint training method and train our vision and natural language projection models for only 1 epoch with a batch size of 256 and a learning rate of $10^{-4}$.
Figure 5: Cross-modality entailment probability of $P_{\text{cross}}(x_{i}^{vision},x_{i}^{NL})=0.5\cdot P(\Omega_{i}^{vision}|\Omega_{i}^{NL})+0.5\cdot P(\Omega_{i}^{NL}|\Omega_{i}^{vision})$ over joint training steps. It can be observed that projection models of vision modality $f_{vision}$ and natural language modality $f_{NL}$ can quickly learn to produce overlapping projections for the same object in the concept space. Such quick convergence allows easy incorporation of new modalities/modalities into the proposed learning system. This joint training takes significantly less time and uses fewer GPU resources than the following BLIP and CLIP models.
Figure 5 illustrates the fast convergence of the proposed projection models on learning to produce overlapping representations of the same objects in the transparent concept space. This joint training also takes significantly less time and uses fewer GPU resources than the following BLIP and CLIP models.
D.2 BLIP
@@ D.5 FLAVA
Appendix E Computation Resources
We run our experiments on a virtual machine (VM) hosted by Microsoft’s Azure. This VM has four NVIDIA A100 PCIe GPUs with 320 GB of total memory.

### v2

@@ C.2 Dataset Preprocessing
Appendix D Model Details
D.1 Backbone Architectures
D.1.1 Vision Modality
ViT. A Vision Transformer Dosovitskiy et al. (2020) pretrained on ImageNet-21k (vit-base-patch16-224) is used as a backbone. The pooled embedding at the [CLS] token is passed to either the projection head (our framework) or the MLP head (baseline).
ResNets. A ResNet-50 He et al. (2015) pretrained on ImageNet-21k is used as a backbone. Its final fully connected layer is replaced by either our projection head or the baseline MLP head.
D.1.2 Natural Language Modality
BERT. A pretrained BERT-Base encoder Devlin et al. (2018) is used as the backbone for the natural-language modality. Analogous to the ViT setup, the pooled representation at the [CLS] position is used as input to both the projection head and the MLP head.
D.2 Baseline MLP Heads
To provide consistent and comparable baselines for our projection models, we attach a simple three-layer MLP head to each backbone, whereas our projection heads consist of only a single linear layer. Unless otherwise specified, the two intermediate layers contain 128 units each. For ResNet-based models, however, the baseline MLP head uses 512 and 256 units in the intermediate layers to accommodate the higher dimensionality of ResNet’s feature representation.
MLP Head Output Dimension.
The output dimension of each baseline MLP is set to the total number of concepts present in the dataset:
$\mathrm{dim(out)}=|\mathcal{Y}_{\text{category}}|+|\mathcal{Y}_{\text{attribute}}|.$
Final predictions are obtained by applying an argmax over the category neurons and a threshold over the attribute neurons.
D.3 Projection Heads
As shown in Algo. 1, our projection head is implemented using two parallel linear layers. Conceptually, the projection head is a single linear mapping from the backbone feature vector into the knowledge space $\mathcal{K}$. For computational convenience, however, we split the backbone feature vector into two equal-sized chunks and apply an independent linear layer to each chunk. This is equivalent to applying one linear layer to the full feature vector, but allows us to separately obtain $(\omega_{\mathrm{min}},\omega_{\Delta})\in\mathcal{K}$ in our framework.
D.4 Training Details
@@
Appendix E Image-Text Matching Experiment Details
E.1 BLIP